# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ C.5 Data-Constrained Scaling Law Fitting Results for EEG
The most notable difference in the EEG scaling results appears in the model-size exponent. For EEG, we obtain an unusually large $\alpha=3.29$, whereas the iEEG experiments produced a much smaller and value of $\alpha=0.3773$. Such a steep exponent suggests that, within the range of model sizes we explored, performance changed very little with additional capacity. A plausible explanation is the behavior of the XXL model: its loss is higher than that of the smaller models, probably due to overfitting. Similar behavior has been reported in very large models trained on fixed data in the Data-constrained scaling law literature(data_constrained_scaling_muennighoff2023scaling). In addition, we observed that losses often rise around epochs 4–8; because our analysis used the midpoint checkpoint of the 32-epoch runs, this choice may also have distorted the fit.
@@ C.6 Extended EEG Downstream Scaling Results
Pretraining Fraction FACED PhysioNet-MI
ACC kappa F1 ACC kappa F1
1% $111\pm 0.000 0.0$$00\pm 0.000 0.0$$036\pm 0.000$ $0.581\pm 0.007$ $0.441\pm 0.009$ $0.582\pm 0.006$
3% $0.467\pm 0.012$ $0.400\pm 0.013$ $0.474\pm 0.012$ $0.624\pm 0.005$ $0.498\pm 0.007$ $0.626\pm 0.005$
9% $0.451\pm 0.017$ $0.380\pm 0.020$ $0.456\pm 0.017$ $0.600\pm 0.005$ $0.467\pm 0.007$ $0.601\pm 0.005$
10% $0.509\pm 0.013$ $0.448\pm 0.014$ $0.520\pm 0.012$ $0.640\pm 0.006$ $0.519\pm 0.009$ $0.641\pm 0.006$
24% $0.525\pm 0.007$ $0.464\pm 0.008$ $0.529\pm 0.006$ $\mathbf{0.650\pm 0.004}$ $\mathbf{0.533\pm 0.005}$ $\mathbf{0.651\pm 0.004}$
50% $0.493\pm 0.007$ $0.426\pm 0.007$ $0.491\pm 0.006$ $0.628\pm 0.006$ $0.504\pm 0.008$ $0.628\pm 0.006$
100% $\mathbf{0.568\pm 0.011}$ $\mathbf{0.513\pm 0.013}$ $\mathbf{0.579\pm 0.009}$ $0.647\pm 0.006$ $0.529\pm 0.008$ $0.649\pm 0.006$
@@
$d_{\text{model}}$ (Params) FACED PhysioNet-MI
ACC kappa F1 ACC kappa F1
256 (13M) $0.488\pm 0.006$ $0.421\pm 0.006$ $0.485\pm 0.005$ $0.630\pm 0.006$ $0.506\pm 0.008$ $0.631\pm 0.007$
512 (51M) $\mathbf{0.568\pm 0.011}$ $\mathbf{0.513\pm 0.013}$ $\mathbf{0.579\pm 0.009}$ $0.647\pm 0.006$ $0.529\pm 0.008$ $0.649\pm 0.006$
768 (114M) $0.477\pm 0.006$ $0.411\pm 0.006$ $0.477\pm 0.006$ $0.659\pm 0.005$ $0.546\pm 0.007$ $0.660\pm 0.005$
1024 (204M) $0.524\pm 0.005$ $0.462\pm 0.005$ $0.523\pm 0.005$ $\mathbf{0.691\pm 0.006}$ $\mathbf{0.588\pm 0.008}$ $\mathbf{0.692\pm 0.006}$
2048 (813M) $0.525\pm 0.012$ $0.462\pm 0.012$ $0.524\pm 0.010$ $0.669\pm 0.011$ $0.560\pm 0.014$ $0.672\pm 0.010$
@@
$d_{\text{model}}$ Epochs FACED PhysioNet-MI
ACC kappa F1 ACC kappa F1
256 2 $153\pm 0.007 0.0$$050\pm 0.007$ $0.103\pm 0.015$ $0.580\pm 0.006$ $0.440\pm 0.007$ $0.582\pm 0.006$
4 $0.463\pm 0.013$ $0.392\pm 0.015$ $0.461\pm 0.014$ $0.604\pm 0.008$ $0.472\pm 0.011$ $0.605\pm 0.008$
8 $0.484\pm 0.005$ $0.417\pm 0.006$ $0.484\pm 0.006$ $0.614\pm 0.003$ $0.485\pm 0.004$ $0.614\pm 0.003$
16 $0.447\pm 0.003$ $0.375\pm 0.003$ $0.443\pm 0.002$ $0.609\pm 0.004$ $0.479\pm 0.005$ $0.610\pm 0.004$
32 $0.488\pm 0.006$ $0.421\pm 0.006$ $0.485\pm 0.005$ $0.615\pm 0.003$ $0.487\pm 0.005$ $0.615\pm 0.004$
512 2 $0.510\pm 0.012$ $0.448\pm 0.013$ $0.517\pm 0.012$ $0.624\pm 0.005$ $0.498\pm 0.007$ $0.625\pm 0.005$
4 $0.539\pm 0.012$ $0.479\pm 0.013$ $0.543\pm 0.010$ $0.654\pm 0.005$ $0.538\pm 0.007$ $0.655\pm 0.005$
8 $0.445\pm 0.008$ $0.373\pm 0.008$ $0.444\pm 0.008$ $0.619\pm 0.004$ $0.491\pm 0.005$ $0.619\pm 0.004$
16 $0.476\pm 0.004$ $0.410\pm 0.004$ $0.477\pm 0.004$ $0.668\pm 0.002$ $0.557\pm 0.003$ $0.669\pm 0.002$
32 $0.493\pm 0.005$ $0.429\pm 0.005$ $0.495\pm 0.005$ $0.680\pm 0.004$ $0.573\pm 0.006$ $0.681\pm 0.004$
768 2 $0.387\pm 0.024$ $0.309\pm 0.027$ $0.386\pm 0.025$ $0.605\pm 0.004$ $0.473\pm 0.005$ $0.606\pm 0.003$
4 $0.511\pm 0.009$ $0.448\pm 0.010$ $0.516\pm 0.010$ $0.636\pm 0.005$ $0.515\pm 0.006$ $0.638\pm 0.005$
8 $0.531\pm 0.004$ $0.470\pm 0.005$ $0.534\pm 0.005$ $0.663\pm 0.004$ $0.550\pm 0.006$ $0.664\pm 0.004$
16 $0.519\pm 0.011$ $0.458\pm 0.012$ $0.518\pm 0.012$ $0.682\pm 0.004$ $0.576\pm 0.005$ $0.683\pm 0.004$
32 $0.526\pm 0.005$ $0.467\pm 0.006$ $0.527\pm 0.005$ $\mathbf{0.690\pm 0.002}$ $\mathbf{0.587\pm 0.002}$ $\mathbf{0.691\pm 0.002}$
1024 2 $0.510\pm 0.010$ $0.449\pm 0.010$ $0.518\pm 0.008$ $0.638\pm 0.005$ $0.518\pm 0.007$ $0.640\pm 0.005$
4 $\mathbf{0.544\pm 0.011}$ $\mathbf{0.486\pm 0.013}$ $\mathbf{0.554\pm 0.010}$ $0.656\pm 0.004$ $0.542\pm 0.005$ $0.658\pm 0.004$
8 $0.462\pm 0.012$ $0.393\pm 0.013$ $0.464\pm 0.012$ $0.591\pm 0.006$ $0.454\pm 0.008$ $0.592\pm 0.007$
16 $0.461\pm 0.007$ $0.390\pm 0.008$ $0.458\pm 0.007$ $0.630\pm 0.008$ $0.507\pm 0.011$ $0.632\pm 0.008$
32 $0.451\pm 0.030$ $0.380\pm 0.033$ $0.449\pm 0.030$ $0.644\pm 0.016$ $0.526\pm 0.022$ $0.646\pm 0.016$
2048 2 $0.380\pm 0.220$ $0.301\pm 0.246$ $0.355\pm 0.261$ $0.620\pm 0.016$ $0.495\pm 0.014$ $0.624\pm 0.009$
4 $111\pm 0.000 0.0$$00\pm 0.000 0.0$$036\pm 0.000$ $0.512\pm 0.131$ $0.351\pm 0.176$ $0.485\pm 0.193$
8 $0.450\pm 0.018$ $0.379\pm 0.021$ $0.450\pm 0.019$ $0.620\pm 0.006$ $0.495\pm 0.007$ $0.623\pm 0.005$
16 $0.514\pm 0.008$ $0.449\pm 0.008$ $0.512\pm 0.007$ $0.658\pm 0.004$ $0.547\pm 0.005$ $0.662\pm 0.004$
32 $0.531\pm 0.014$ $0.470\pm 0.015$ $0.514\pm 0.013$ $0.661\pm 0.009$ $0.550\pm 0.012$ $0.665\pm 0.009$
@@ D.1 Comprehensive DIVER Downstream Task Results
We present the detailed numerical values. The EEG performance table is provided in Table 14. The iEEG performance table is provided in Table 12 and Table 13. For Neuroprobe binary-label (iEEG task), we evaluated DIVERTiny/I/0.1s (with $d_{\text{model}}=256$ and patch size 0.1s, pretrained on iEEG for 32 epochs) in both frozen (linear probing, red) and fine-tuned (blue) configurations against Linear Laplacian STFT, BrainBERT and PopT baselines on Figure 2. Table 13 presents comprehensive results comparing DIVER with baseline models across all 15 Neuroprobe tasks. For Neuroprobe multi-label, our model are still overperform the linear baseline. Results are reported as mean AUROC $\pm$ SEM across subjects, trials, and cross-validation folds. DIVER consistently outperformed baseline models across the majority of tasks in both evaluation settings. The model demonstrated particularly strong performance on language-related tasks (speech decoding, word prediction, onset detection) and auditory tasks (volume, pitch), with finetuning providing additional gains over frozen features. These results validate that self-supervised pretraining on iEEG data produces representations that transfer effectively to diverse downstream neural decoding tasks. In the multi-label setting, DIVER again outperformed the linear baseline for both frozen and fully fine-tuned models. [^4] A multi-label option is available in the most recent release of the Neuroprobe code, but it has not yet been documented in the paper. Consequently, we compare only against the linear (STFT–Laplacian) baseline, witch can be easily compute by running their released code.
@@
Models Overall Sentence Onset Speech Volume
Linear Laplacian STFT $0.660\pm 0.005$ $0.891\pm 0.018$ $0.883\pm 0.018$ $0.717\pm 0.032$
BrainBERT (frozen) $0.586\pm 0.004$ $0.757\pm 0.027$ $0.611\pm 0.022$ $0.583\pm 0.010$
PopT $0.545\pm 0.006$ $0.689\pm 0.050$ $0.677\pm 0.044$ $0.576\pm 0.018$
DIVERTiny/I/0.1s $0.662\pm 0.008$ $0.924\pm 0.009$ $\mathbf{0.900\pm 0.011}$ $0.699\pm 0.020$
DIVERTiny/I/0.1s(frozen) $\mathbf{0.676\pm 0.007}$ $\mathbf{0.930\pm 0.008}$ $0.896\pm 0.012$ $\mathbf{0.717\pm 0.018}$
@@ D.2 Performance Evaluation across DIVER-1 model configurations
Tiny/I/0.1s (frozen) $0.896\pm 0.012$ $\textbf{0.930}\pm\textbf{0.008}$ $\textbf{0.717}\pm\textbf{0.018}$ $\textbf{0.589}\pm\textbf{0.007}$ $\textbf{0.676}\pm\textbf{0.007}$}
@@
And for MAYO (seizure detection), we chosed only 1 sec model, because MAYO has 6 s window that is too long for 0.1 s model’s context. We compared the frozen and full finetuning model in each model size in Table 16. In contrast to the Neuroprobe results, we found that full fine-tuning outperformed the frozen for the Tiny, Small and XL model, even though it was slightly worse for the Large and Base models. Since our preprocessing clips signal amplitudes above a certain threshold (200 μV), so pretrained-dataset’s distribution can differ from seizure data, which contains spikes with much larger amplitudes; under this distribution shift, fully fine-tuned models may therefore tend to achieve better performance. For the linear baseline, the highest performance was obtained with the Large model, whereas under full-finetuning the Tiny model achieved the best performance.
@@ D.3 Ablations
For EEG downstream tasks, detailed results are described in Table 18. The ablation of encoder components yielded a significant performance decline in both EEG tasks, indicating that the encoder components are crucial contributors to overall performance. Removing the multi-domain reconstruction task, STCPE and spectral feature embedding resulted in a notable performance degradation as well. Ablation of channel-wise patch embedding components induced inconsistent results across tasks; for FACED, performance dropped while for PhysioNet-MI, performance slightly improved.
@@
Dataset ablations and dataset size-based comparisons The other iEEG models (PopT (chau2025popt) and BrainBERT (wang2023brainbert)) are pretrained on the BrainTreebank (BTB) (wang2024brain) datasets—the precursor to Neuroprobe. Therefore, we compare downstream performance when we use BTB exclusively (Figure 12). We trained DIVERTiny/I/0.1s models with BTB and size-variations of our self-collected iEEG datasets. For the model trained on our self-collected data of the same size as BTB, the linear-probing results were lower than those trained with BTB only. However, when we increased the size of the self-collected dataset (approximately ×16 and ×64 size of BTB), performance surpassed the BTB-only setting. This shows that, with sufficient data, it is possible to achieve higher performance even if the distribution of the pretraining dataset differs from that of the downstream tasks.
@@
Table 19: Model comparison with a shared pretraining dataset. Each row indicates a model architecture with its pretraining dataset in parentheses. Models pretrained on the same BTB dataset are compared to isolate architectural effects. The results are reported as mean AUROC ± SEM across multiple subjects, trials and folds. The DIVER model was pretrained for 32 epochs.
speech onset
BrainBERT (frozen) ((2024neuroprobe) $0.611\pm 0.022$ $0.757\pm 0.027$
BrainBERT (our training code) $0.575\pm 0.018$ $0.659\pm 0.026$
PopT (our training code) $0.702\pm 0.029$ $0.780\pm 0.025$
PopT ((2024neuroprobe) $0.677\pm 0.044$ $0.689\pm 0.050$
DIVERTiny/I/0.1s (frozen) $\mathbf{0.770\pm 0.028}$ $\mathbf{0.859\pm 0.018}$

### v2

@@ C.5 Data-Constrained Scaling Law Fitting Results for EEG
The most notable difference in the EEG scaling results appears in the model-size exponent. For EEG, we obtain an unusually large $\alpha=3.29$, whereas the iEEG experiments produced a much smaller and value of $\alpha=0.3773$. Such a steep exponent suggests that, within the range of model sizes we explored, performance changed very little with additional capacity. A plausible explanation is the behavior of the XXL model: its loss is higher than that of the smaller models, probably due to overfitting. Similar behavior has been reported in very large models trained on fixed data in the Data-constrained scaling law literature(Muennighoff et al., 2023). In addition, we observed that losses often rise around epochs 4–8; because our analysis used the midpoint checkpoint of the 32-epoch runs, this choice may also have distorted the fit.
@@ C.6 Extended EEG Downstream Scaling Results
[…]
black Pretraining Fraction FACED PhysioNet-MI ACC kappa F1 ACC kappa F1 1% $111\pm 0.000 0.0$$00\pm 0.000 0.0$$036\pm 0.000$ $0.581\pm 0.007$ $0.441\pm 0.009$ $0.582\pm 0.006$ 3% $0.467\pm 0.012$ $0.400\pm 0.013$ $0.474\pm 0.012$ $0.624\pm 0.005$ $0.498\pm 0.007$ $0.626\pm 0.005$ 9% $0.451\pm 0.017$ $0.380\pm 0.020$ $0.456\pm 0.017$ $0.600\pm 0.005$ $0.467\pm 0.007$ $0.601\pm 0.005$ 10% $0.509\pm 0.013$ $0.448\pm 0.014$ $0.520\pm 0.012$ $0.640\pm 0.006$ $0.519\pm 0.009$ $0.641\pm 0.006$ 24% $0.525\pm 0.007$ $0.464\pm 0.008$ $0.529\pm 0.006$ $\mathbf{0.650\pm 0.004}$ $\mathbf{0.533\pm 0.005}$ $\mathbf{0.651\pm 0.004}$ 50% $0.493\pm 0.007$ $0.426\pm 0.007$ $0.491\pm 0.006$ $0.628\pm 0.006$ $0.504\pm 0.008$ $0.628\pm 0.006$ 100% $\mathbf{0.568\pm 0.011}$ $\mathbf{0.513\pm 0.013}$ $\mathbf{0.579\pm 0.009}$ $0.647\pm 0.006$ $0.529\pm 0.008$ $0.649\pm 0.006$
@@
[…]
black $d_{\text{model}}$ (Params) FACED PhysioNet-MI ACC kappa F1 ACC kappa F1 256 (13M) $0.488\pm 0.006$ $0.421\pm 0.006$ $0.485\pm 0.005$ $0.630\pm 0.006$ $0.506\pm 0.008$ $0.631\pm 0.007$ 512 (51M) $\mathbf{0.568\pm 0.011}$ $\mathbf{0.513\pm 0.013}$ $\mathbf{0.579\pm 0.009}$ $0.647\pm 0.006$ $0.529\pm 0.008$ $0.649\pm 0.006$ 768 (114M) $0.477\pm 0.006$ $0.411\pm 0.006$ $0.477\pm 0.006$ $0.659\pm 0.005$ $0.546\pm 0.007$ $0.660\pm 0.005$ 1024 (204M) $0.524\pm 0.005$ $0.462\pm 0.005$ $0.523\pm 0.005$ $\mathbf{0.691\pm 0.006}$ $\mathbf{0.588\pm 0.008}$ $\mathbf{0.692\pm 0.006}$ 2048 (813M) $0.525\pm 0.012$ $0.462\pm 0.012$ $0.524\pm 0.010$ $0.669\pm 0.011$ $0.560\pm 0.014$ $0.672\pm 0.010$
@@
[…]
black $d_{\text{model}}$ Epochs FACED PhysioNet-MI ACC kappa F1 ACC kappa F1 256 2 $153\pm 0.007 0.0$$050\pm 0.007$ $0.103\pm 0.015$ $0.580\pm 0.006$ $0.440\pm 0.007$ $0.582\pm 0.006$ 4 $0.463\pm 0.013$ $0.392\pm 0.015$ $0.461\pm 0.014$ $0.604\pm 0.008$ $0.472\pm 0.011$ $0.605\pm 0.008$ 8 $0.484\pm 0.005$ $0.417\pm 0.006$ $0.484\pm 0.006$ $0.614\pm 0.003$ $0.485\pm 0.004$ $0.614\pm 0.003$ 16 $0.447\pm 0.003$ $0.375\pm 0.003$ $0.443\pm 0.002$ $0.609\pm 0.004$ $0.479\pm 0.005$ $0.610\pm 0.004$ 32 $0.488\pm 0.006$ $0.421\pm 0.006$ $0.485\pm 0.005$ $0.615\pm 0.003$ $0.487\pm 0.005$ $0.615\pm 0.004$ 512 2 $0.510\pm 0.012$ $0.448\pm 0.013$ $0.517\pm 0.012$ $0.624\pm 0.005$ $0.498\pm 0.007$ $0.625\pm 0.005$ 4 $0.539\pm 0.012$ $0.479\pm 0.013$ $0.543\pm 0.010$ $0.654\pm 0.005$ $0.538\pm 0.007$ $0.655\pm 0.005$ 8 $0.445\pm 0.008$ $0.373\pm 0.008$ $0.444\pm 0.008$ $0.619\pm 0.004$ $0.491\pm 0.005$ $0.619\pm 0.004$ 16 $0.476\pm 0.004$ $0.410\pm 0.004$ $0.477\pm 0.004$ $0.668\pm 0.002$ $0.557\pm 0.003$ $0.669\pm 0.002$ 32 $0.493\pm 0.005$ $0.429\pm 0.005$ $0.495\pm 0.005$ $0.680\pm 0.004$ $0.573\pm 0.006$ $0.681\pm 0.004$ 768 2 $0.387\pm 0.024$ $0.309\pm 0.027$ $0.386\pm 0.025$ $0.605\pm 0.004$ $0.473\pm 0.005$ $0.606\pm 0.003$ 4 $0.511\pm 0.009$ $0.448\pm 0.010$ $0.516\pm 0.010$ $0.636\pm 0.005$ $0.515\pm 0.006$ $0.638\pm 0.005$ 8 $0.531\pm 0.004$ $0.470\pm 0.005$ $0.534\pm 0.005$ $0.663\pm 0.004$ $0.550\pm 0.006$ $0.664\pm 0.004$ 16 $0.519\pm 0.011$ $0.458\pm 0.012$ $0.518\pm 0.012$ $0.682\pm 0.004$ $0.576\pm 0.005$ $0.683\pm 0.004$ 32 $0.526\pm 0.005$ $0.467\pm 0.006$ $0.527\pm 0.005$ $\mathbf{0.690\pm 0.002}$ $\mathbf{0.587\pm 0.002}$ $\mathbf{0.691\pm 0.002}$ 1024 2 $0.510\pm 0.010$ $0.449\pm 0.010$ $0.518\pm 0.008$ $0.638\pm 0.005$ $0.518\pm 0.007$ $0.640\pm 0.005$ 4 $\mathbf{0.544\pm 0.011}$ $\mathbf{0.486\pm 0.013}$ $\mathbf{0.554\pm 0.010}$ $0.656\pm 0.004$ $0.542\pm 0.005$ $0.658\pm 0.004$ 8 $0.462\pm 0.012$ $0.393\pm 0.013$ $0.464\pm 0.012$ $0.591\pm 0.006$ $0.454\pm 0.008$ $0.592\pm 0.007$ 16 $0.461\pm 0.007$ $0.390\pm 0.008$ $0.458\pm 0.007$ $0.630\pm 0.008$ $0.507\pm 0.011$ $0.632\pm 0.008$ 32 $0.451\pm 0.030$ $0.380\pm 0.033$ $0.449\pm 0.030$ $0.644\pm 0.016$ $0.526\pm 0.022$ $0.646\pm 0.016$ 2048 2 $0.380\pm 0.220$ $0.301\pm 0.246$ $0.355\pm 0.261$ $0.620\pm 0.016$ $0.495\pm 0.014$ $0.624\pm 0.009$ 4 $111\pm 0.000 0.0$$00\pm 0.000 0.0$$036\pm 0.000$ $0.512\pm 0.131$ $0.351\pm 0.176$ $0.485\pm 0.193$ 8 $0.450\pm 0.018$ $0.379\pm 0.021$ $0.450\pm 0.019$ $0.620\pm 0.006$ $0.495\pm 0.007$ $0.623\pm 0.005$ 16 $0.514\pm 0.008$ $0.449\pm 0.008$ $0.512\pm 0.007$ $0.658\pm 0.004$ $0.547\pm 0.005$ $0.662\pm 0.004$ 32 $0.531\pm 0.014$ $0.470\pm 0.015$ $0.514\pm 0.013$ $0.661\pm 0.009$ $0.550\pm 0.012$ $0.665\pm 0.009$
@@ D.1 Comprehensive DIVER Downstream Task Results
We present the detailed numerical values. The EEG performance table is provided in Table 15. The iEEG performance table is provided in Table 13 and Table 14. For Neuroprobe binary-label (iEEG task), we evaluated DIVERTiny/I/0.1s (with $d_{\text{model}}=256$ and patch size 0.1s, pretrained on iEEG for 32 epochs) in both frozen (linear probing, red) and fine-tuned (black) configurations against Linear Laplacian STFT, BrainBERT and PopT baselines on Figure 3. Table 14 presents comprehensive results comparing DIVER with baseline models across all 15 Neuroprobe tasks. For Neuroprobe multi-label, our model are still overperform the linear baseline. Results are reported as mean AUROC $\pm$ SEM across subjects, trials, and cross-validation folds. DIVER consistently outperformed baseline models across the majority of tasks in both evaluation settings. The model demonstrated particularly strong performance on language-related tasks (speech decoding, word prediction, onset detection) and auditory tasks (volume, pitch), with finetuning providing additional gains over frozen features. These results validate that self-supervised pretraining on iEEG data produces representations that transfer effectively to diverse downstream neural decoding tasks. In the multi-label setting, DIVER again outperformed the linear baseline for both frozen and fully fine-tuned models. [^3] A multi-label option is available in the most recent release of the Neuroprobe code, but it has not yet been documented in the paper. Consequently, we compare only against the linear (STFT–Laplacian) baseline, witch can be easily compute by running their released code.
@@
[…]
black Models Overall Sentence Onset Speech Volume Linear Laplacian STFT $0.660\pm 0.005$ $0.891\pm 0.018$ $0.883\pm 0.018$ $0.717\pm 0.032$ BrainBERT (frozen) $0.586\pm 0.004$ $0.757\pm 0.027$ $0.611\pm 0.022$ $0.583\pm 0.010$ PopT $0.545\pm 0.006$ $0.689\pm 0.050$ $0.677\pm 0.044$ $0.576\pm 0.018$ DIVERTiny/I/0.1s $0.662\pm 0.008$ $0.924\pm 0.009$ $\mathbf{0.900\pm 0.011}$ $0.699\pm 0.020$ DIVERTiny/I/0.1s(frozen) $\mathbf{0.676\pm 0.007}$ $\mathbf{0.930\pm 0.008}$ $0.896\pm 0.012$ $\mathbf{0.717\pm 0.018}$
@@ D.2 Performance Evaluation across DIVER-1 Model configurations
Tiny/I/0.1s (frozen) $0.896\pm 0.012$ $\textbf{0.930}\pm\textbf{0.008}$ $\textbf{0.717}\pm\textbf{0.018}$ $\textbf{0.589}\pm\textbf{0.007}$ $\color[rgb]{0,0,0}\textbf{0.676}\pm\textbf{0.007}$}
@@
[…]
@@ D.3 Ablations
For EEG downstream tasks, detailed results are described in Table 19. The ablation of encoder components yielded a significant performance decline in both EEG tasks, indicating that the encoder components are crucial contributors to overall performance. Removing the multi-domain reconstruction task, STCPE and spectral feature embedding resulted in a notable performance degradation as well. Ablation of the channel-wise patch embedding components led to inconsistent effects across tasks: performance dropped on FACED, whereas performance on PhysioNet-MI did not degrade meaningfully.
@@
Dataset ablations and dataset size-based comparisons The other iEEG models (PopT (Chau et al., 2025) and BrainBERT (Wang et al., 2023)) are pretrained on the BrainTreebank (BTB) (Wang et al., 2024a) datasets—the precursor to Neuroprobe. Therefore, we compare downstream performance when we use BTB exclusively (Figure 13). We trained DIVERTiny/I/0.1s models with BTB and size-variations of our self-collected iEEG datasets. For the model trained on our self-collected data of the same size as BTB, the linear-probing results were lower than those trained with BTB only. However, when we increased the size of the self-collected dataset (approximately ×16 and ×64 size of BTB), performance surpassed the BTB-only setting. This shows that, with sufficient data, it is possible to achieve higher performance even if the distribution of the pretraining dataset differs from that of the downstream tasks.